OpenStudy (mariebrittany07):

how do you figure out the problem 5x-1=x+15

OpenStudy (anonymous):

first subtract x from the right side and subtract it from the right. So it's 4x-1=15. Then add one to both sides so its 4x=16. Divide both sides by 4. x=4

OpenStudy (anonymous):

sorry *subtract x from left and right

OpenStudy (mariebrittany07):

so I subtract x from left to right so it is like this 5x-1=x+15-x

OpenStudy (mariebrittany07):

Am I right or wrong

OpenStudy (anonymous):

Well usually you just right it underneath. So like 5x-1=x+15 -x -x ------------- 4x-1=-15 +1 +1 4x=16

OpenStudy (anonymous):

then divide both my 4. and get x=4

OpenStudy (mariebrittany07):

oh okay i get it now so i was almost right

OpenStudy (mariebrittany07):

do i 4x/4 = 16/4